A movie theater has two ticket prices: $8 for adults and $5 for children. I f the box office took in $4116 from the sale of 600 tickets how many tickerts of each kind were sold?
Answer by ikleyn(52802)

.
 a +  c =  600,   (1)    (counting tickets)
8a + 5c = 4116.   (2)    (counting dollars)


To solve, express c = 600-a from (1) and substitute it into (2). You will get the single equation for a:

8a + 5*(600-a) = 4116.

Simplify and solve it.  Then evaluate  c = 600-a.


It is very standard problem on tickets (actually, on solving systems of linear equations in 2 unknowns).

See the lesson
    - Using systems of equations to solve problems on tickets
in this site.